OpenStudy (anonymous):

Given 10 rooms and 5 people, ask each person to choose one room to stay. What is the possibility that no people select same room?

OpenStudy (anonymous):

i think its 1/10 * 1/9* 1/8 * 1/7* 1/6 though probability is not my strong point a second opinion is essential!

OpenStudy (anonymous):

Thanks. So the possibility is very little based on your equation. I was thinking it is P(10,5)/10^5=0.3

OpenStudy (phi):

The number of ways to assign rooms, allowing multiple people to choose the same room is 10^5 the number of ways to assign rooms so only 1 person per room is 10*9*8*7*6 this ratio is your answer.

OpenStudy (anonymous):

Same here. While then what is the possibility that there are two people select one room? C(10,1)*C(9,3)/10^5?

OpenStudy (phi):

I think it's 10*(9*8*7) for the numerator. i.e. P(9,3) not C(9,3)

OpenStudy (phi):

2nd opinion welcome!

OpenStudy (anonymous):

Agree. P(9,3) instead of C(9,3)

OpenStudy (anonymous):

identical to "birthday problem" with 10 days and 5 people.

OpenStudy (anonymous):

pretty pretty sure it is \[\frac{10P5}{10^{10}}\] but i would not swear to it

OpenStudy (anonymous):

Ooh, confusing. I'll try...

OpenStudy (anonymous):

well that is wrong. try \[\frac{10\times 9\times 8\times 7\times 6}{10^5}\]

OpenStudy (anonymous):

you can read the "birthday problem" here http://en.wikipedia.org/wiki/Birthday_problem#Understanding_the_problem then change 365 to 10 and 23 to 5

OpenStudy (anonymous):

I think it is kind of birthday problem. While if ask this: Given these 10 rooms and 5 people, what is the (expected) number of rooms have two or more people in it?

OpenStudy (anonymous):

first person picks a room. he has 10 choice. then second person. now he has 9 choices etc

OpenStudy (anonymous):

to the number of ways they can all pick different rooms is \[10\times 9\times 8\times 7\times 6\times 5\]

OpenStudy (anonymous):

and the number of way total that 5 people can pick from 10 rooms is \[10^5\]

OpenStudy (anonymous):

so pretty sure that my second answer (not the first) is right

OpenStudy (anonymous):

For the chance of no same room, It is P(10,5)/10^5=0.3, isn't it?

OpenStudy (anonymous):

oh and since phi said exactly the same thing i am really pretty sure

OpenStudy (anonymous):

@phi not sure why you stopped the numerator at 6 though. fifth person needs to pick a different room as well

OpenStudy (anonymous):

oops now i see why! sorry

OpenStudy (phi):

I'm sure about the first problem. But I'm also sure I have a bug in the 2nd question, the prob of two people picking the same room.

OpenStudy (anonymous):

i only see one question. hmmm

OpenStudy (anonymous):

Second question is : what is the possibility that there are two people select one room? C(10,1)*P(9,3)/10^5?

OpenStudy (anonymous):

Third one is :): Given these 10 rooms and 5 people, what is the (expected) number of rooms have two or more people in it?

OpenStudy (phi):

On the 3rd question, prob of multiple occupancy is 1- prob(unique rooms) = 1-0.3=0.7

OpenStudy (anonymous):

0.7 is the prob of there is people share one room. the question is , how many rooms we can expect to see that are shared by more than one people

OpenStudy (phi):

Back to the 2nd q. I think we need a factor of 5*4 to account for the different pairs of people, so the numerator is 10*9*8*7*5*4

OpenStudy (anonymous):

phi: Agree. For second question it is 10*9*8*7*5*4/10^5

OpenStudy (phi):

Do you know?

OpenStudy (anonymous):

Now I am little bit sure about the second question Just no idea for the third one

OpenStudy (anonymous):

question #2 is 1 - question 1

OpenStudy (anonymous):

unless i messed that up too!

OpenStudy (phi):

I interpreted it as exactly 2 which is different from not 1

OpenStudy (anonymous):

oh yes. exactly one room has two people in it is not what i said for sure

OpenStudy (phi):

Q3: how about this? 1 * (.7 * 0.3^4) + 2*(.7^2 * 0.3^3) + 3*(0.7^3 * 0.3^2)+ 4*(0.7^4 * 0.3)+5*(0.7^5)

OpenStudy (phi):

= 1.25

OpenStudy (anonymous):

let me be quiet that has nothing to do with it sorry

OpenStudy (phi):

expected value is a weighted sum: number of rooms * probability of event occurring

OpenStudy (anonymous):

I am trying to understand your equation phi

OpenStudy (phi):

It was meant as an idea.... I do not believe it is correct.

OpenStudy (phi):

Bad news on 2nd q. 10*9*8*7*5*4/10^5 = 1.008. So this is definitely wrong.

OpenStudy (anonymous):

Now I am thinking the second question is C(5,2)*C(10,1)*P(9,3)/10^5=0.72, right?

OpenStudy (anonymous):

Sorry, it is 0.504 based on previous equation

OpenStudy (phi):

Q3: expected # of shared rooms = 0.8146 1 1 1 1 1 0.3024 2 1 1 1 0.5040 2 2 1 0.1080 3 1 1 0.0720 3 2 0.0090 4 1 0.0045 5 0.0001 -------- 1.0000 expected # shared = 1*0.5040 + 2*0.108 + 1*0.072 + 2*0.009 + 0.0045 + 0.0001

OpenStudy (anonymous):

Thank you Phi!